OpenStudy (anonymous):

solve the systems of equation: 2w - 5z = 13 6w + 3z = 10 use any method.

OpenStudy (amistre64):

you made it, you solve it ;)

OpenStudy (anonymous):

what?

OpenStudy (amistre64):

id timesicate the top by -3 maybe

OpenStudy (amistre64):

timesiply? multiplicate? .... somethin

hero (hero):

amistre, speak English

OpenStudy (amistre64):

my english speakatorial is on the fritz and jam i think

hero (hero):

Have you thought about retiring? I think your English speaking abilities are starting to decline, lol

OpenStudy (amistre64):

i had to read "a brave new world" and "mans search for meaning" so I had to forget a few things to make them fit ;)

OpenStudy (amistre64):

orish, what methods do you feel comfortable with?

OpenStudy (anonymous):

any

OpenStudy (amistre64):

good, then how far have you gotten :)

OpenStudy (amistre64):

2w - 5z = 13 ; *-3 6w + 3z = 10 -6w +15z = -39 6w + 3z = 10 --------------- 18z = -29 z = -29/18 and simplify as wanted

OpenStudy (amistre64):

2w - 5z = 13 6w + 3z = 10 ; *5/3 2w - 5z = 13(3)/3 10w +5z = 50/3 --------------- 12w = 89/3 w = 89/36 ... and simplify

OpenStudy (anonymous):

thats fast :)

OpenStudy (amistre64):

and efficient lol